JET-COOLED LASER-INDUCED FLUORESCENCE SPECTROSCOPY OF LARGE SECONDARY ALKOXY RADICALS

\underline{JINJUN LIU}, MING-WEI CHEN, TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, the Ohio State University, 120 W. 18^{th} Ave., Columbus, Ohio 43210; W. L. MEERTS, Radboud University, Institute for Molecules and Materials, Heyendaalseweg 135, NL-6525 AJ Nijmegen, The Netherlands.

The B̃ ← X̃ laser-induced fluorescence (LIF) spectra of jet-cooled 2-pentoxy and 2-hexoxy have been recorded. The observed rotational and fine structure of the strongest vibronic bands has been simulated with ab initio calculated rotational constants for both the X̃ and B̃ states, as well as the electron spin-rotation constants of the X̃ state and the transition dipole moments, which are predicted based on the transferability of these quantities in an “orbital-fixed coordinate system” using iso-propoxy as the reference molecule. It is suggested by ab initio calculations that the lowest two electronic (X̃ and Ã) states of secondary alkoxy radicals have a small energy separation on the order of 100 cm^{−1}. The energy ordering of these two nearly degenerate states has been determined by comparing the experimentally determined rotational constants and the transition dipole moments to the predicted ones. Molecular constants derived in fitting the rotational and fine structure of the experimental spectra using an evolutionary algorithm (EA) enabled unambiguous assignment of the observed vibronic bands to different conformers of 2-pentoxy and 2-hexoxy. Based on the results of these two radicals, the strongest vibronic bands in the LIF spectra of larger secondary alkoxies were also assigned.